OpenStudy (anonymous):

if f and g are increasing functions on an interval I in R, show that f + g is an increasing function on I. If fi s also strictly increasing on I then f + g is strictly increasing on I. @Number Theory

OpenStudy (anonymous):

can we simply say f(x1)< = f(x2) < =.....<=f(xn) and g(x1) <=g(x2)<=.....<=g(xn) and f(x1) + g(x1) <=f(x2)+g(x2) <=.....<=f(xn)+g(xn) and (f + g)(x1) <= (f+g)(x2)<=....<= (f+g)(xn) so (f+g)(x) is also an increasing function